A boy whirls a storm in a horizontal circle of radius 1.5 and a Height 2 meter above level ground. The string breaks, and the stone flies off horizontally and Strikes the ground after traveling horizontally a distance of 10 meter. What is the magnitude of the centripetal acceleration of the stone during the circular motion?
1
Expert's answer
2019-12-09T09:03:52-0500

Time of flight:


"t=\sqrt{\frac{2h}{g}}=\sqrt{\frac{2(2)}{9.8}}=0.639\ s"

"v=\frac{d}{t}=\frac{10}{0.639}=15.6\frac{m}{s}"

The magnitude of the centripetal acceleration of the stone during the circular motion:


"a=\frac{v^2}{r}=\frac{15.6^2}{1.5}=160\frac{m}{s^2}"